OpenStudy (kanwal32):

The first term of a convex polygon are in AP. The smallest angle is 120 degree and the common difference is 5 degree find the number of sides of the polygon

OpenStudy (kanwal32):

i used the approach that a polygon cannot have an angle 180 degree cause then the polygon will not close i took the last term as 175 degree and used the formula of ap answer is comming 12 but the answer in the testbook is given 9

OpenStudy (kanwal32):

@ganeshie8

OpenStudy (kanwal32):

@Kainui

OpenStudy (kanwal32):

@paki @ParthKohli

OpenStudy (kanwal32):

@mathslover

OpenStudy (kanwal32):

@hartnn

OpenStudy (kanwal32):

@ganeshie8

OpenStudy (kanwal32):

@Kainui

OpenStudy (er.mohd.amir):

two ans are 16 and 9 u find wrong second one since T16=120+(16-1)*5=120+75=195 and angle are not more then 180 so ans is 9 is right.

OpenStudy (kanwal32):

@dan815